Which of the following statements is (are) correct?
(x) An externality is the uncompensated impact of one person’s actions on the well-being of a bystander.
(y) An example of an externality that causes market failure is the case of a copper refinery that does not bear the entire cost of the smoke it emits.
(z) If an externality, such as air or water pollution, is present in a market, economic efficiency may be enhanced by government intervention that reduces the presence of the pollution.
A. (x), (y) and (z)
B. (x) and (y) only
C. (x) and (z) only
D. (y) and (z) only
E. (z) only

Causes of market failure include
A. market power and incorrect forecasts of consumer demand.
B. externalities and market power.
C. externalities and foreign competition.
D. incorrect forecasts of consumer demand and foreign competition.
E. Both A and D are correct


9. A typical worker in Italy can produce 24 units of product in an 8 hour day, where a typical worker in Poland produces 25 units of product in a 10 hour day. We can conclude that
(x) worker productivity in Poland is higher than in Italy.
(y) the standard of living will be higher in Italy than in Poland.
(z) productivity for the Polish worker is 2 1/2 units per hour and 3 units per hour for the Italian worker.
A. (x), (y) and (z)
B. (x) and (y) only
C. (x) and (z) only
D. (y) and (z) only
E. (y) only
Answer #1
  1. Option a) is correct as all the three statements are true. X and y are self-explanatory. Z is true as the negative effects of the externality vanishes.
  2. Option b) is correct.

9)    Worker productivity in Italy = 24 / 8 = 3 and in Poland = 25/10 = 2.5. Z is true. X is false. Option d) is true.